# KTester: Leveraging Domain and Testing Knowledge for More Effective LLM-based Test Generation

### v1

@@ Abstract.
Automated unit test generation using large language models (LLMs) holds great promise but often struggles with generating tests that are both correct and maintainable in real-world projects. This paper presents KTester, a novel framework that integrates project-specific knowledge and testing domain knowledge to enhance LLM-based test generation. Our approach first extracts project structure and usage knowledge through static analysis, which provides rich context for the model. It then employs a testing-domain-knowledge-guided separation of test case design and test method generation, combined with a multi-perspective prompting strategy that guides the LLM to consider diverse testing heuristics. The generated tests follow structured templates, improving clarity and maintainability. We evaluate KTester on multiple open-source projects, comparing it against state-of-the-art LLM-based baselines using automatic correctness and coverage metrics, as well as a human study assessing readability and maintainability. Results demonstrate that KTester significantly outperforms existing methods across six key metrics, improving execution pass rate by 5.69% and line coverage by 8.83% over the strongest baseline, while requiring less time and generating fewer test cases. Human evaluators also rate the tests produced by KTester significantly higher in terms of correctness, readability, and maintainability, confirming the practical advantages of our knowledge-driven framework.
@@ 1. Introduction
Unit testing is a fundamental practice in software development that verifies whether a method behaves as expected, playing a key role in ensuring software correctness, maintainability, and enabling regression testing (leung1989insights). As the first line of defense in the development lifecycle, it helps detect and localize bugs early, preventing their propagation (runeson2006survey). However, writing high-quality unit tests manually is often time-consuming and error-prone, as it requires developers to deeply understand the method’s behavior, construct valid input states, and define precise assertions (daka2014survey).
For a method under test (i.e., often called as the focal method), its well-constructed unit test typically consists of two parts: the test prefix, which sets up the necessary objects and environment to bring the system into a testable state, and the test oracle, which verifies that the actual output matches the expected behavior (DBLP:journals/tse/BarrHMSY15). High-quality unit tests not only improve software reliability but also serve as living documentation, enabling easier code comprehension and maintenance (fraser2014large).
To reduce manual effort, automated test generation techniques have been proposed to generate a suite of unit tests with the main goal of maximizing the coverage in the software under test. Traditional unit test generation techniques include search-based (harman2009theoretical; blasi2022call; delgado2022interevo), constraint-based methods (ernst2007daikon; csallner2008dysy; xiao2013characteristic), and random-based strategies (zeller2019fuzzing; pacheco2007feedback). While effective in achieving code coverage, the generated tests are often hard to read and maintain, limiting practical adoption (almasi2017industrial).
@@
As illustrated in Figure 1 to 3, these issues commonly arise due to the absence of project-specific knowledge and test domain expertise in existing LLM-based test generation methods. For instance, Figure 1 shows a test that incorrectly constructs the labels Map used by SparkApplication, inserting the key “dragName” instead of the correct constant Constants.DAG_NAME_LABE. This error arises from the test’s lack of awareness of constructor dependencies and configuration requirements. Figure 2 depicts a test with insufficient assert statements, weakening its effectiveness. Meanwhile, Figure 3 shows tests that are inconsistently structured, relying on hard-coded values and missing essential setup logic, which results in fragile and difficult-to-maintain code.
@@
Our approach’s core innovations are twofold. First, the explicit incorporation of project-aware knowledge provides the LLM with rich, contextualized information about the codebase, enabling generation of more accurate and meaningful tests. Second, we introduce a testing domain knowledge–guided separation of test design and test generation and multi-perspective test case design prompting strategy. By decoupling the “what to test” from “how to test”, we make the testing intent explicit and improve the semantic clarity and maintainability of the generated test code.
We validate our approach on several real-world open-source projects, comparing it against state-of-the-art LLM-based baselines using both automated metrics (e.g., execution pass rate, line coverage) and a human study assessing readability, maintainability, and test intent clarity. Our findings show that (1) our method consistently outperforms baselines across six key metrics, improving execution pass rate by 5.69% and line coverage by 8.83% compared to the strongest baseline, while takes shorter time and generate less test cases; (2) the modular test case transformation component, which embodies the testing-principle-guided design, has the largest individual impact, with removal causing a drop of 24.08% in execution pass rate and 12.61% in line coverage; and (3) human evaluators rate our generated tests significantly higher in correctness, readability, and maintainability, confirming the practical advantages of our knowledge-driven framework.
@@
All data/code used in this study is provided in the package (Replication_Package).
@@ 2. Approach
  • Project Knowledge Awareness: We statically analyze the entire codebase to extract structured information about classes, methods, usage patterns, dependencies, and documentation. This knowledge is later used to resolve object instantiation, locate related APIs, and reduce hallucinations during code generation.
  • Testing Knowledge Awareness: We explicitly guide the LLM to design test cases from multiple testing perspectives (e.g., control flow, boundary, exception handling), instead of treating the task as a black-box code-to-code translation.
  • Modular Generation Pipeline: We separate test class framework construction, scenario-specific test design, and final test case synthesis into distinct stages, improving modularity, reusability, and interpretability of generated tests.
@@ 2.1.2. Project Usage Knowledge Mining
To mine project usage knowledge, we begin by identifying the relevant caller context and then extract semantically meaningful execution traces that illustrate how the focal method is invoked in practice.
Path-sensitive Usage Trace Extraction. For each discovered caller method, we construct a Control Flow Graph (CFG) (allen1970control) and perform intra-procedural slicing to extract execution paths from the method’s entry point to the focal method invocation. By pruning unrelated branches and removing irrelevant operations, we isolate minimal yet semantically rich usage traces. These traces, which demonstrate how the focal class is instantiated and how arguments are prepared, are incorporated into the generation prompt as part of the Project Usage Knowledge. This provides the LLM with grounded, context-aware examples that help it synthesize test inputs aligned with the project’s actual usage conventions.
@@ 2.1.3. Function-Level Index Construction
To enable scalable and consistent test generation, each method or constructor is transformed into a knowledge unit containing its signature, documentation, and dependencies (invoked methods and accessed fields). These units are stored in an indexable format, forming a function-level knowledge base that supports efficient retrieval of related methods, usage contexts, and class-level information during prompt construction.
Unlike ad hoc, per-method extraction, our analysis is conducted once per project and reused across focal methods. The knowledge base is also incrementally updatable—only changed entities need to be reprocessed—making it practical for real-world, evolving codebases. While our method is language-agnostic, we focus on Java due to its popularity and mature analysis tooling(javaparser; spoon_2025). We leverage Spoon (spoon_2025) to extract ASTs, control flow, and dependency graphs efficiently.
@@ 2.2.1. Test Class Framework Generation
The first step prompts the LLM to generate a test class framework that establishes the necessary environment for unit testing. Unlike prior methods (chen_chatunitest_2024; wang_hits_2024) that directly prompt the LLM with a focal method and expect complete test cases in one shot, KTester decouples the generation process by first constructing a reusable, project-aware test class framework. This decision is grounded in a key insight: test generation is not merely a method-level translation task, but a context-sensitive activity requiring awareness of surrounding project structure, usage conventions, and domain-specific testing patterns. As illustrated in Figure 7, we design a prompt template grounded in project knowledge and test domain knowledge to guide this process. The prompt is composed of several key components, including core task and instructions, focal method and class, relevant project knowledge, test class template, and output specification.
@@
Figure 7. Prompt for test class framework generation.
@@ 2.2.2. Test Cases Design
KTester adopts a multi-view guidance strategy grounded in established software testing principles. The LLM is explicitly instructed to design test cases from multiple perspectives—such as functional behavior, boundary conditions, and exception handling—and to organize them into logical groups based on shared testing intent. Each test case is represented internally in a machine-readable intermediate format that records structured natural-language descriptions of the scenario. In our implementation, this intermediate representation is realized using a lightweight JSON structure, and an illustrative example is provided in our replication package (Replication_Package). This design intentionally decouples test intent from concrete executable code, improving clarity and maintainability. The intermediate representations are subsequently translated into framework-specific test code.
@@
Figure 8. Prompts for multi-view test case design.
@@ 2.2.3. Test Method Transformation
Figure 9. Prompt for test method transformation.
Usage Context Retrieval. To enhance generation quality, we retrieve methods that are functionally related to the focal method and its dependencies, offering a broader context of real-world usage patterns. Based on our observations, related methods are identified based on shared usage of functions and fields. We define the similarity between two functions $a$ and $b$ using a Jaccard-based metric (Jaccard1902) by combining with method usage similarity $Sim_{m}(a,b)$ and field usage similarity $Sim_{f}(a,b)$:
@@ 2.2.4. Test Class Integration
  • Test Method Deduplication. For test methods and general helper methods, we identify duplicates by comparing their method names. When two methods share the same name, we further compare their bodies and retain the version with the longer implementation, using length as a proxy for logical richness and completeness. This deduplication addresses a common pattern where the LLM initially produces simple placeholder tests and later generates more complete versions with the same or similar names, ensuring that the final output preserves the more comprehensive implementation.
  • Non-conflicting Insertion. For all other non-conflicting methods, we append them directly to the test class. This modular addition supports test diversity and extensibility without disrupting the original structure.
The integration process leverages standard AST-based static analysis tools to parse and manipulate the Java code structure safely and consistently. This automation ensures that the final test class is executable, maintainable, and free of duplicate or conflicting code.
@@ 2.2.5. Test Class Refinement
In this stage, KTester validates the generated test class by checking for compilation and execution errors. Despite the rich contextual information used during generation, LLM outputs may still contain issues such as hallucinated import statements, improper mocking, incorrect usage of private methods, or runtime errors during test execution. To address these problems, we adopt an iterative refinement process consisting of both static validation and dynamic repair.
@@
LLM-based Repair. For errors arising from semantic misunderstandings or subtle logic issues that cannot be fixed by static rules alone, KTester uses LLM-based repair. This involves extracting detailed compile-time error messages, runtime error logs, and failing test feedback. This information, combined with the focal method context and relevant function signatures, is incorporated into a repair prompt that is fed back to the LLM. The model then performs targeted correction to resolve issues such as incorrect assertions, missing exception handling, or logic bugs causing runtime failures.
@@ 3.1.1. Dataset
To evaluate the effectiveness of KTester, we adopt the HITS dataset (wang_hits_2024), which is constructed from 10 popular open-source Java projects spanning various domains (e.g., microservices, command-line tools, event engines). This dataset specifically targets complex methods—defined as those with cyclomatic complexity greater than 10—making it well-suited for assessing how effectively LLMs handle methods with intricate control flow and dependencies. The dataset comprises 110 tasks, each consisting of a focal method and its containing class.
@@
3.1.2. Baselines
ChatUnitTest (chen_chatunitest_2024) provides the focal method and handcrafted context to the LLM, and applies iterative repair using error feedback from failed executions.
ChatTester (yuan2023no) is similar to ChatUnitTest but differs in how it constructs the focal method’s context, building it incrementally as needed.
HITS (wang_hits_2024) decomposes the focal method into slices and prompts the LLM to generate test cases for each slice, which are then merged into a complete test suite. This fine-grained approach helps mitigate the difficulty LLMs face when reasoning about complex logic holistically.
For fair comparison, we adapt official or open-source implementations of these methods (zju-chatunitest_2025). ChatUnitTest and ChatTester generate one test class per focal method, while HITS’s slice-level tests are merged into a single class to ensure consistent execution and setup. All baselines and KTester use the same LLM backend—gpt-4o-mini (chatgpt_model)—with default API settings with temperature=1.0, ensuring consistency in generation quality, efficiency, and cost.
The SBST–LLM hybrid method, UTGen (deljouyi2025leveraging), employs a traditional SBST tool to generate initial test cases, then leverages an LLM to rename functions and variables in order to improve test readability. Since UTGen rely on an LLM’s ability on code understanding instead of enhancing code coverage, we follow the initial configuration provided in its replication package (UTGen) (EvoSuite (fraser_evosuite_2011) as the generation tool and CodeLlama-7b (codellama) used for refinement) when conducting our experiments to evaluate its performance.

### v2

# Knowledge Matters: Injecting Project and Testing Knowledge into LLM-based Unit Test Generation

@@ Abstract.
Automated unit test generation using large language models (LLMs) holds great promise but often struggles with generating tests that are both correct and maintainable in real-world projects. This paper presents KTester, a novel framework that integrates project-specific knowledge and testing domain knowledge to enhance LLM-based test generation. Our approach first extracts project structure and usage knowledge through static analysis, which provides rich context for the model. It then employs a testing-domain-knowledge-guided separation of test case design and test method generation, combined with a multi-perspective prompting strategy that guides the LLM to consider diverse testing heuristics. The generated tests follow structured templates, improving clarity and maintainability. We evaluate KTester on multiple open-source projects, comparing it against state-of-the-art LLM-based baselines using automatic correctness and coverage metrics, as well as a human study assessing readability and maintainability. Results demonstrate that KTester significantly outperforms existing methods across six key metrics, improving execution pass rate by 5.03% and line coverage by 11.67% over the strongest baseline, while requiring less time and generating fewer test cases. Human evaluators also rate the tests produced by KTester significantly higher in terms of correctness, readability, and maintainability, confirming the practical advantages of our knowledge-driven framework.
Unit Test, Large Language Models, Domain Knowledge
*M. Liu is the corresponding author. A. Li, M. Liu, Z. Chen, P. Zheng, Z. Li, D. Dai, Y. Wang and Z. Zheng are with the School of Software Engineering and the Zhuhai Key Laboratory of Trusted Large Language Models, Sun Yat-sen University, Zhuhai, China. Email: lianj8@mail2.sysu.edu.cn, liumw26@mail.sysu.edu.cn, {chenzhx236, peizh3, lizk8,daidk}@mail2.sysu.edu.cn, {wangylin36, zhzibin}@mail.sysu.edu.cn.
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3787769isbn: 979-8-4007-2025-3/2026/04ccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Software development process management
@@ 1. Introduction
Unit testing is a fundamental practice in software development that verifies whether a method behaves as expected, playing a key role in ensuring software correctness, maintainability, and enabling regression testing (Leung and White, 1989). As the first line of defense in the development lifecycle, it helps detect and localize bugs early, preventing their propagation (Runeson, 2006). However, writing high-quality unit tests manually is often time-consuming and error-prone, as it requires developers to deeply understand the method’s behavior, construct valid input states, and define precise assertions (Daka and Fraser, 2014).
For a method under test (i.e., often called the focal method), a well-constructed unit test usually has two parts: the test prefix, which prepares the objects and environment needed to bring the system into a testable state, and the test oracle, which verifies that the actual output matches the expected behavior (Barr et al., 2015). High-quality unit tests improve software reliability and also serve as living documentation, facilitating code comprehension and maintenance (Fraser and Arcuri, 2014).
To reduce manual effort, automated test generation techniques have been proposed to generate a suite of unit tests with the main goal of maximizing the coverage in the software under test. Traditional unit test generation techniques include search-based (Harman and McMinn, 2009; Blasi et al., 2023; Delgado-Pérez et al., 2023), constraint-based methods (Ernst et al., 2007; Csallner et al., 2008; Xiao et al., 2013), and random-based strategies (Zeller et al., 2019; Pacheco et al., 2007). While effective in achieving code coverage, the generated tests are often hard to read and maintain, limiting practical adoption (Almasi et al., 2017).
Motivational Examples A Motivational Examples(a): Setting Tested Object Incorrectly
@@
As illustrated in Figure 1-3, these issues commonly stem from the lack of project-specific knowledge and test domain expertise in existing LLM-based test generation methods. For instance, Figure 1 shows an incorrect construction of the labels Map used by SparkApplication, inserting the key “dragName” instead of the correct constant Constants.DAG_NAME_LABE. This error arises from a lack of awareness of constructor dependencies and configuration requirements. Figure 2 depicts a test with insufficient assert statements, weakening its effectiveness. Meanwhile, Figure 3 shows an inconsistent test structure with hard-coded values and missing essential setup logic, which results in fragile and difficult-to-maintain code.
@@
Motivation Examples B Motivational Examples(b): Insufficient Assert Statements
Our approach’s core innovations are twofold. First, the explicit incorporation of project-aware knowledge provides the LLM with rich, contextualized information about the codebase, leading to more accurate and meaningful tests. Second, we introduce a testing domain knowledge–guided separation of test design and test generation and multi-perspective test case design prompting strategy. By decoupling the “what to test” from “how to test”, we make the testing intent explicit and improve the semantic clarity and maintainability of the generated test code.
We validate our approach on several real-world open-source projects, comparing it against state-of-the-art LLM-based baselines using both automated metrics (e.g., execution pass rate, line coverage) and a human study assessing readability, maintainability, and test intent clarity. Our findings show that (1) our method consistently outperforms baselines across six key metrics, improving execution pass rate by 5.03% and line coverage by 11.67% compared to the strongest baseline, while takes shorter time and generate less test cases; (2) the modular test case transformation component, which embodies the testing-principle-guided design, has the largest individual impact, with removal causing a drop of 11.45% in execution pass rate and 13.39% in line coverage; and (3) human evaluators rate our generated tests significantly higher in correctness, readability, and maintainability, confirming the practical advantages of our knowledge-driven framework.
@@
Motivational Examples(c) Motivational Examples(c): Hard Encoding Values
@@
All data/code used in this study is provided in the package (36).
@@ 2. Approach
Project Knowledge Awareness. We statically analyze the entire codebase to extract structured information about classes, methods, usage patterns, dependencies, and documentation. This knowledge is later used to resolve object instantiation, locate related APIs, and reduce hallucinations during code generation.
Figure 4. Test Class Example
Test Class Example The framework of test class, includes classes, methods, usage patterns, dependencies, and documentation.
Testing Knowledge Awareness. We explicitly guide the LLM to design test cases from multiple testing perspectives (e.g., control flow, boundary, exception handling), instead of treating the task as a black-box code-to-code translation.
Modular Generation Pipeline. We separate test class framework construction, scenario-specific test design, and final test case synthesis into distinct stages, improving modularity, reusability, and interpretability of generated tests.
@@ 2.1.2. Project Usage Knowledge Mining
To mine project usage knowledge, we first identify the relevant caller context and then extract semantically meaningful execution traces showing how the focal method is invoked in practice.
Path-sensitive Usage Trace Extraction. For each discovered caller method, we construct a Control Flow Graph (CFG) (Allen, 1970) and perform intra-procedural slicing to extract execution paths from the method’s entry point to the focal method invocation. By pruning unrelated branches and removing irrelevant operations, we isolate minimal yet semantically rich usage traces. These traces, which demonstrate how the focal class is instantiated and how arguments are prepared, are incorporated into the generation prompt as part of the Project Usage Knowledge. This provides the LLM with grounded, context-aware examples that help it synthesize test inputs aligned with the project’s actual usage conventions.
@@
Invocation Examples Illustration of extracting invocation examples from a caller method via control flow pruning.
@@ 2.1.3. Function-Level Index Construction
To enable scalable and consistent test generation, each method or constructor is transformed into a knowledge unit with its signature, documentation, and dependencies (invoked methods and accessed fields). These units are stored in an indexable format, forming a function-level knowledge base that enables efficient retrieval of related methods, usage contexts, and class-level information during prompt construction.
Unlike ad hoc, per-method extraction, our analysis is conducted once per project and reused across focal methods. The knowledge base is also incrementally updatable, requiring reprocessing only changed entities, which makes it practical for evolving codebases. While our method is language-agnostic, we focus on Java due to its popularity and mature analysis tooling(29; 32). We use Spoon (32) to efficiently extract ASTs, control flow, and dependency graphs.
@@ 2.2. Test Class Generation
Figure 7. Prompt for test class framework generation.
Prompt for test class framework generation. Prompt for test class framework generation.
@@ 2.2.1. Test Class Framework Generation
The first step prompts the LLM to generate a test class framework that establishes the necessary environment for unit testing. Unlike prior methods (Chen et al., 2024; Wang et al., 2024) that directly prompt the LLM with a focal method and expect complete test cases in one shot, KTester decouples the generation process by first constructing a reusable, project-aware test class framework. This decision is grounded in a key insight: test generation is not merely a method-level translation task, but a context-sensitive activity requiring awareness of surrounding project structure, usage conventions, and domain-specific testing patterns. As illustrated in Figure 7, we design a prompt template grounded in project knowledge and test domain knowledge to guide this process. The prompt is composed of several key components, including core task and instructions, focal method and class, relevant project knowledge, test class template, and output specification.
@@
Figure 8. Prompts for multi-view test case design.
Prompts for multi-view test case design. Prompts for multi-view test case design.
@@ 2.2.2. Test Cases Design
KTester adopts a multi-view guidance strategy grounded in established software testing principles. The LLM is explicitly instructed to design test cases from multiple perspectives—such as functional behavior, boundary conditions, and exception handling—and to organize them into logical groups based on shared testing intent. Each test case is represented internally in a machine-readable intermediate format that records structured natural-language descriptions of the scenario. In our implementation, this intermediate representation is realized using a lightweight JSON structure, and an illustrative example is provided in our replication package (36). This design intentionally decouples test intent from concrete executable code, improving clarity and maintainability. The intermediate representations are subsequently translated into framework-specific test code.
@@ 2.2.3. Test Method Transformation
Usage Context Retrieval. To enhance generation quality, we retrieve methods that are functionally related to the focal method and its dependencies, offering a broader context of real-world usage patterns. Based on our observations, related methods are identified based on shared usage of functions and fields. We define the similarity between two functions $a$ and $b$ using a Jaccard-based metric (Jaccard, 1902) by combining with method usage similarity $Sim_{m}(a,b)$ and field usage similarity $Sim_{f}(a,b)$:
@@ 2.2.4. Test Class Integration
Test Method Deduplication. For test methods and general helper methods, we identify duplicates by method name. When two methods share the same name, we further compare their bodies and retain the longer implementation as a proxy for logical richness. This deduplication addresses a common pattern where the LLM first generates simple placeholder tests and later generates more complete versions with the same or similar names, ensuring that the final output preserves the more comprehensive implementation.
Non-conflicting Insertion. For all other non-conflicting methods, we append them directly to the test class. This modular addition supports test diversity and extensibility without disrupting the original structure. The integration process leverages standard AST-based static analysis tools to parse and manipulate the Java code structure safely and consistently. This automation ensures that the final test class is executable, maintainable, and free of duplicate or conflicting code.
@@
Figure 9. Prompt for test method trasformation.
Prompt for test method trasformation.
@@ 2.2.5. Test Class Refinement
In this stage, KTester validates the generated test class by checking compilation and execution errors. Despite the rich contextual guidance, LLM outputs may still contain issues such as hallucinated import statements, improper mocking, incorrect usage of private methods, or runtime errors during test execution. To address these problems, we adopt an iterative refinement process with static validation and dynamic repair.
@@
LLM-based Repair. For errors arising from semantic misunderstandings or subtle logic issues that cannot be fixed by static rules, KTester applies LLM-based repair using compile-time errors, runtime logs, and failing test feedback. This information, combined with the focal method context and relevant function signatures, is incorporated into a repair prompt that is fed back to the LLM. The model then performs targeted correction of issues such as incorrect assertions, missing exception handling, or logic bugs causing runtime failures.
@@ 3.1.1. Dataset
To evaluate the effectiveness of KTester, we adopt the HITS dataset (Wang et al., 2024), which is constructed from 10 popular open-source Java projects spanning various domains (e.g., microservices, command-line tools, event engines). This dataset specifically targets complex methods—defined as those with cyclomatic complexity greater than 10—making it well-suited for assessing how effectively LLMs handle methods with intricate control flow and dependencies. The dataset comprises 110 tasks, each consisting of a focal method and its containing class.
3.1.2. Baselines
ChatUnitTest (Chen et al., 2024) provides the focal method and handcrafted context to the LLM, and applies iterative repair using error feedback from failed executions.
ChatTester (Yuan et al., 2024) is similar to ChatUnitTest but differs in how it constructs the focal method’s context, building it incrementally as needed.
HITS (Wang et al., 2024) decomposes the focal method into slices and prompts the LLM to generate test cases for each slice, which are then merged into a complete test suite. This fine-grained approach helps mitigate the difficulty LLMs face when reasoning about complex logic holistically.
For fair comparison, we adapt official or open-source implementations of these methods (30). ChatUnitTest and ChatTester generate one test class per focal method, while HITS’s slice-level tests are merged into a single class to ensure consistent execution and setup. All baselines and KTester use the same LLM backend—gpt-4o-mini (31)—with temperature=0.5, ensuring consistency in generation quality, efficiency, and cost.
The SBST–LLM hybrid method, UTGen (Deljouyi et al., 2025), employs a traditional SBST tool to generate initial test cases, then leverages an LLM to rename functions and variables in order to improve test readability. Since UTGen rely on an LLM’s ability on code understanding instead of enhancing code coverage, we follow the initial configuration provided in its replication package (51) (EvoSuite (Fraser and Arcuri, 2011) as the generation tool and CodeLlama-7b (12) used for refinement) when conducting our experiments to evaluate its performance.